# On codes induced from Hadamard matrices 111 Keywords: Hadamard, code, linear, convolutional, self-dual, dual-containing, quantum code, linear complementary AMS Classification 2020: 94B05, 94B10, 15B99

### v1

@@ 2 Linear block and convolutional codes induced from Hadamard matrices
Let $B=H[11..24][1..24]$. Then both $A$ and $B$ generate $[24,12,7]_{5}$ codes over $GF(5)$. Let $G(z)=A+iBz$ where $i=\sqrt{-1}=2$ in $GF(5)$. Then by Proposition 2.3, $G(z)$ generates a convolutional self-dual $[24,12,12;1,14)$ code over $GF(5)$. From this a QECC convolutional code of length $24$ and distance $14$ is derived over $GF(5)$.
@@
Let $H=H(40),A=H[1..20][1..40],B=H[21..40][1..40]$. Over $GF(5)$ $H$ has rank $20$ but $A$ has rank $10$ over $GF(5)$. Now $C=H[1..10][1..40]$ has rank $10$ over $GF(5)$ and generates a $[40,10,16]$ LCD code over $GF(5)$. Over $GF(3)$ define $G(z)=A+iBz$ where $i=\sqrt{-1}$ in $GF(3^{2})$. Then $G(z)$ generates a self-dual convolutional $(40,20,20;1,d)$ code where $d=d(A)+d(B)$, Proposition 2.3. From this a QECC convolutional code of length $40$ and distance $d$ is obtained.
@@
$H=H(36),A=H[1..18][1..36],B=H[19..36][1..36]$. Over $GF(3)$ $A$ generates a self-dual $[36,18,12]_{3}$ code as does $B$. Define $G(z)=A+iBz$ where $i=\sqrt{-1}=2$ in $GF(5)$ and then $G(z)$ generates a convolutional self-dual code $(36,18,18;1,d)$ where $d=d(A)+d(B)$ from which a QECC convolutional code of length $36$ and distance $d$ is obtained.
@@ 2.1 Ternary codes from Hadamard matrices
Lemma 2.7 shows that when $3\mid n$ then $\operatorname{rank}(H)\leq\frac{n}{2}$ in $\mathbb{Z}_{3}=GF(3)$ for a Hadamard matrix of size $n$. For a Paley-Hadamard matrix $H$ of size $n$, $\operatorname{rank}(H)=\frac{n}{2}$. Is it true in other cases? In cases where the rank is $\frac{n}{2}$ a self-dual $[n,\frac{n}{2},d]_{3}$ code is constructible from any $\frac{n}{2}$ independent rows of the Hadamard matrix.
@@
Let $H$ be a Hadamard matrix of size $n$ such that $3\mid n$ and that $\operatorname{rank}(H)=\frac{n}{2}$. Then any submatrix of size $\frac{n}{2}\times n$ of rank $\frac{n}{2}$ over $GF(3)$ generates a self-dual $[n,\frac{n}{2},d]_{3}$ code.
@@
Let $H=H(12k)$ be a Hadamard matrix of rank $6k$ over $GF(3)$. Then $6k$ linearly independent rows over $GF(3)$ of $H$ generate a self-dual $[n,\frac{n}{2},d]_{3}$ code.
It is interesting to find the distances attained. For a self-dual $[n,\frac{n}{2},d]_{3}$ ternary code it is known that $d\leq\lfloor\frac{n}{12}\rfloor+3$ [20]. For $n=12k$ extremal ternary self-dual codes exist for lengths $n=12,24,36,48,60$ and do not exist for $n=72,96,120$ and for $n\geq 144$. Now by Corollary 2.17 $\frac{n}{2}$ linearly independent rows over $GF(3)$ of a Hadamard matrix of size $n=12k$ generate a self-dual ternary code; when $n=12,24,36,48,60$ it is verified by computer that these are optimal.
@@
In characteristic $3$ the non-zero coefficients in a sum of rows are ±$\pm 1$ only. Interchanging rows of a Hadamard matrix or multiplying any row by $-1$ results in a Hadamard matrix. Thus taking the relevant rows and placing them in the first $r$ places and multiplying the row by $-1$ if the coefficient is $-1$ results in a Hadamard matrix whose sum of the first $r$ rows is the same as the vector sum of the required rows. ∎
Thus if a lower bound on the support of the sum of the first $s$ rows of a particular type of Hadamard matrix over $\mathbb{Z}_{3}$ can be obtained then distances of the unit-derived codes from such a matrix $H$ are calculated and also the distances of the self-dual codes $(I,\alpha H)$ as in Proposition 2.2 are obtained.
The following Proposition is a special case of Proposition 2.1.
@@
Let $H$ be a Hadamard $n\times n$ and $3\nmid n$. Then any $r$ rows of $H$ generates an LCD ternary $[n,r,d]_{3}$ code over $\mathbb{Z}_{3}$.
Application: With $n=20$ the following LCD codes are obtained: $[20,5,10]_{3},[20,6,10]_{3},[20,10,6]_{3},[20,11,5]_{3},[20,13,4]_{3}$ With $n=28$ the following LCD codes are obtained: $[28,7,12],[28,14,6],[28,18,4]$.
@@
(ii) Suppose $A$ consists of $r$ rows of $H$ with $r>\frac{n}{2}$ and $B$ consists of the other $(n-r)$ rows of $H$. Define $B_{1}=\left(\begin{smallmatrix}\underline{0}\\ B\end{smallmatrix}\right)$ where $\underline{0}$ is the zero $(2n-r)\times n$ matrix. Then $G(z)=A+iB_{1}x$ generates a dual-containing convolutional $(n,r,n-r;1,d)$ code in $GF(3^{2})$ where $i=\sqrt{-1}$.
Proposition 2.5 can be used to calculate or estimate the distance $d$ in part (ii) of Proposition 2.20.

### v2

@@ 2 Linear block and convolutional codes induced from Hadamard matrices
Let $B=H[11..24][1..24]$. Then both $A$ and $B$ generate $24,12,7]_{5}$ codes over $GF(5)$. Let $G(z)=A+iBz$ where $i=\sqrt{-1}=2$ in $GF(5)$. Then by Proposition 2.3, $G(z)$ generates a convolutional self-dual $[24,12,12;1,14)$ code over $GF(5)$. From this a QECC convolutional code of length $24$ and distance $14$ is derived over $GF(5)$.
@@
Let H=H(40),A=H[1..20][1..40],B=H[21..40][1..40]H=H(40),A=H[1..20][1..40],B=H[21..40][1..40]. Over $GF(5)$ $H$ has rank $20$ but $A$ has rank $10$ over $GF(5)$. Now $C=H[1..10][1..40]$ has rank $10$ over $GF(5)$ and generates a $[40,10,16]$ LCD code over $GF(5)$. Over $GF(3)$ define $G(z)=A+iBz$ where $i=\sqrt{-1}$ in $GF(3^{2})$. Then $G(z)$ generates a self-dual convolutional $(40,20,20;1,d)$ code where $d=d(A)+d(B)$, Proposition 2.3. From this a QECC convolutional code of length $40$ and distance $d$ is obtained.
@@
H=H(36),A=H[1..18][1..36],B=H[19..36][1..36]H=H(36),A=H[1..18][1..36],B=H[19..36][1..36]. Over $GF(3)$ AA generates a self-dual $36,18,12]_{3}$ code as does $B$. Define $G(z)=A+iBz$ where $i=\sqrt{-1}=2$ in $GF(5)$ and then $G(z)$ generates a convolutional self-dual code $(36,18,18;1,d)$ where $d=d(A)+d(B)$ from which a QECC convolutional code of length $36$ and distance $d$ is obtained.
@@ 2.1 Ternary codes from Hadamard matrices
Lemma 2.7 shows that when $3\mid n$ then $\operatorname{rank}(H)\leq\frac{n}{2}$ in $\mathbb{Z}_{3}=GF(3)$ for a Hadamard matrix of size $n$. For a Paley-Hadamard matrix $H$ of size $n$, $\operatorname{rank}(H)=\frac{n}{2}$. Is it true in other cases? In cases where the rank is $\frac{n}{2}$ a self-dual $n,\frac{n}{2},d]_{3}$ code is constructible from any $\frac{n}{2}$ independent rows of the Hadamard matrix.
@@
Let $H$ be a Hadamard matrix of size $n$ such that $3\mid n$ and that $\operatorname{rank}(H)=\frac{n}{2}$. Then any submatrix of size $\frac{n}{2}\times n$ of rank $\frac{n}{2}$ over $GF(3)$ generates a self-dual $n,\frac{n}{2},d]_{3}$ code.
@@
Let $H=H(12k)$ be a Hadamard matrix of rank $6k$ over $GF(3)$. Then $6k$ linearly independent rows over $GF(3)$ of $H$ generate a self-dual $n,\frac{n}{2},d]_{3}$ code.
It is interesting to find the distances attained. For a self-dual $n,\frac{n}{2},d]_{3}$ ternary code it is known that $d\leq\lfloor\frac{n}{12}\rfloor+3$ [29]. For $n=12k$ extremal ternary self-dual codes exist for lengths $n=12,24,36,48,60$ and do not exist for $n=72,96,120$ and for $n\geq 144$. Now by Corollary 2.17 $\frac{n}{2}$ linearly independent rows over $GF(3)$ of a Hadamard matrix of size $n=12k$ generate a self-dual ternary code; when $n=12,24,36,48,60$ it is verified by computer that these are optimal.
@@
In characteristic $3$ the non-zero coefficients in a sum of rows are $\pm 1$ only. Interchanging rows of a Hadamard matrix or multiplying any row by 1-1 results in a Hadamard matrix. Thus taking the relevant rows and placing them in the first $r$ places and multiplying the row by 1-1 if the coefficient is 1-1 results in a Hadamard matrix whose sum of the first $r$ rows is the same as the vector sum of the required rows. ∎
Thus if a lower bound on the support of the sum of the first $s$ rows of a particular type of Hadamard matrix over $\mathbb{Z}_{3}$ can be obtained then distances of the unit-derived codes from such a matrix $H$ are calculated and also the distances of the self-dual codes $(I,\alpha H)$ as in Proposition 2.2 are obtained. The following Proposition is a special case of Proposition 2.1.
@@
Let $H$ be a Hadamard $n\times n$ and $3\nmid n$. Then any $r$ rows of $H$ generates an LCD ternary $n,r,d]_{3}$ code over $\mathbb{Z}_{3}$.
Application: With $n=20$ the following LCD codes are obtained:
$20,5,10]_{3},[20,6,10]_{3},[20,10,6]_{3},[20,11,5]_{3},[20,13,4]_{3}$
With $n=28$ the following LCD codes are obtained: $[28,7,12],[28,14,6],[28,18,4]$.
@@
(ii) Suppose $A$ consists of $r$ rows of $H$ with $r>\frac{n}{2}$ and $B$ consists of the other (nr)(n-r) rows of $H$. Define $B_{1}=\left(\begin{smallmatrix}\underline{0}\\ B\end{smallmatrix}\right)$ where $\underline{0}$ is the zero $(2n-r)\times n$ matrix. Then $G(z)=A+iB_{1}x$ generates a dual-containing convolutional (n,r,nr;1,d)(n,r,n-r;1,d) code in $GF(3^{2})$ where $i=\sqrt{-1}$.
Proposition 2.5 is used to calculate the distance $d$ in part (ii) of Proposition 2.20.